OpenStudy (anonymous):

Solve x = log_3 15. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

x ≈ 2.56 x ≈ 3.12 x ≈ 2.46 x ≈ 3.67

OpenStudy (helder_edwin):

use the same rule as before \[ \large \log_315=\frac{\log 15}{\log 3} \]

OpenStudy (helder_edwin):

use your calculator

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

?

OpenStudy (helder_edwin):

yes. are you using a calculator or you're guessing